OpenStudy (anonymous):

integrating (1/(x^2 - 4x))dx The root of x^2 - 4x is 4, then x^2 - 4x = (x-4)(x-4) = (x-4)^2 So (1/(x^2 - 4x))dx = (1/((x-4)^2)dx Making u = x-4, du/dx=1 and dx=du, then int (1/(x^2 - 4x))dx = int(1/(u^2))du int(1/(u^2))du = -1/u +c = 1/(x+4)+c But this result doen't match with the right one after doing a numerical test. The correct answer is: (1/4)*ln((x-4)/x)) +c. What's wrong in the first way?

OpenStudy (anonymous):

this line is incorrect \[x^2 - 4x = (x-4)(x-4) = (x-4)^2\]

OpenStudy (anonymous):

\[(x-4)^2=x^2-8x+16\] not \(x^2-4x\)

OpenStudy (anonymous):

\(x^2-4x=x(x-4)\) so you can use partial fractions for this one

OpenStudy (anonymous):

Thanks satellite. I reached x2−4x=x(x−4). Could you tell me why I can't use this rule: ax^2+bx+c =a(x-x1)(x-x2) on the first polynomial? I thought it would be right if I made: x2−4x=1*(x-4)(x-4)=(x-4)^2. What's wrong with this? Thanks again.